# Benchmarking neutrino-nucleus quasielastic scattering model predictions against a missing energy profile obtained using a monoenergetic neutrino beam

### v4

@@ I Introduction
Precise modeling of neutrino–nucleus interactions is essential for current and future neutrino oscillation experiments, where systematic uncertainties associated with nuclear effects directly impact the interpretation of measured signals [9]. Current and future generation neutrino oscillation experiments heavily rely on Monte Carlo event generators to predict neutrino interactions in the GeV regime. Until recently, neutrino generators mostly relied on inclusive theoretical model implementations, providing differential cross section predictions only as a function of final-state lepton kinematics. Over the last few years, generators have undergone an overhaul that accommodates the implementation of more exclusive models by relying on the factorization approach [30, 10]. This has notably allowed for the implementation of shell-based models such as the Benhar Spectral Function (SF) [15, 17, 12]. Recently, the unfactorized energy-dependent relativistic mean field (ED-RMF) model was implemented into the NEUT generator using a tabulated hadron tensor approach [36]. These models have been benchmarked against exclusive electron scattering measurements, and their validation using exclusive neutrino scattering measurements has recently begun [25, 24, 23, 39].
@@
A major challenge in measuring neutrino scattering cross sections arises from the broadband nature of accelerator neutrino beams. The wide energy spectrum complicates the reconstruction of the incoming neutrino energy and smears out final-state observables, making it difficult to disentangle the modeling of the nuclear ground state from nuclear effects. An exception to this comes from neutrinos produced by kaon decay at rest (K-DAR). In this process, a kaon undergoes a two-body decay producing a neutrino ($K^{+}\rightarrow\mu^{+}\nu_{\mu}$) with a well defined energy of 235.5 MeV, with a branching ratio of 63.6% [28].
@@
Recently, the JSNS2 collaboration has released its first measurement of the missing energy spectrum using neutrinos from a K-DAR source on a carbon target [35]. This measurement can be used to study the neutron spectral function of carbon, which is a widely used nuclear target in experiments (for example, by experiments such as T2K [1], MINERvA [8], and NOvA [14]). These experiments operate at energies of $\approx 1$ GeV, so K-DAR neutrinos are comparable to the regions of interest for accelerator-based experiments (such as T2K and NOvA, and in the future, DUNE [7] and Hyper-K [2]). Moreover, these energies are particularly relevant for testing the approximations on which neutrino interaction generators are built. For example, many generators rely on the plane wave impulse approximation (PWIA), which assumes that the neutrino scatters off a single isolated nucleon and that the outgoing particles propagate as undistorted plane waves, neglecting correlations and interactions with the residual nucleus. The energy region of the K-DAR neutrinos is low enough that the energy transfers involved in charged-current (CC) interactions become sensitive to nuclear medium effects (for example, related to the nuclear ground state, Pauli blocking, or nucleon-nucleon correlations), which provides a unique opportunity to probe them directly.
@@
In this work, we focus on benchmarking three exclusive shell-based models implemented in the NEUT event generator [30] (version 6.0.3) against the JSNS2 measurement of K-DAR neutrinos. We choose to focus on the NEUT event generator because, at the time of writing, it offers the widest choice of exclusive shell-based models within a single generator, allowing us to test the models in a consistent framework.
@@ II Interaction Models
[…]
@@
where $m_{j}$ and $s$ are the third component of the total angular momentum of the hole state and the spin of the scattered nucleon state, respectively. $j$ denotes the total angular momentum, and $2j+1$ is the total occupancy of a given shell. The hadron tensor $J^{\mu}_{\kappa,m_{j},s}$ is described by single-particle RMF states [25] as
@@
Here, $\Psi_{\kappa}^{m_{j}}(\mathbf{p})$ is the initial-state RMF wavefunction, $\Gamma^{\mu}$ is the transition operator for the CCQE interaction in the CC2 form [33], and $\bar{\psi}_{s}(\mathbf{p}+\mathbf{q},\mathbf{p_{N}})$ is the final-state wavefunction that is the solution to the Dirac equation in the presence of a nuclear potential. $\mathbf{p}$ is the momentum of the initial nucleon, $\mathbf{p_{N}}$ is the momentum of the final-state nucleon, and $\mathbf{q}$ is the three momentum transfer. The model has multiple nuclear potentials available: ED-RMF [25], energy-dependent $A$-independent relativistic optical potentials (EDAI) [21], the real-only component of the EDAI optical potential (rEDAI) and no potential which simplifies to the relativistic plane wave impulse approximation (RPWIA). The inclusion of the nuclear potential in the final state is not present in the SF models and provides additional information about elastic FSI that the scattered nucleon undergoes. In order to model the nucleus in a similar fashion to the SF model, without the need for factorisation, the missing energy profile is modeled using Gaussian nuclear shells [36]; the Gaussian shell structure for carbon is given in Table 1 and a “background” shell contribution is included that is modeled in three distinct regions in $E_{m}$. The shell positions are chosen to be solutions to RMF equations or are taken from recent measurements where they exist. The shell occupancies and widths are chosen to mimic the SF [15, 16] following the procedure described in Ref. [23]. The ED-RMF potential has a natural implementation of Pauli blocking due to the orthogonality of the wavefunctions for the initial and final states [25].
@@
The NEUT FSI cascade is based on the mean free path of a particle travelling through the nuclear medium [30]. Particles are tracked and stepped by a fixed size of 0.2 fm through the nuclear medium until it leaves the nucleus. During this process, nucleon rescattering based on the Bertini model [19], pion absorption based on Ref. [42] can occur, leading creation and ejection of other particles from the nucleus. The NEUT FSI cascade is inherently an inelastic process; during nucleon rescatters, both nucleons involved in the interaction are ejected. In contrast, the ED-RMF model incorporates the final-state potential into calculating the kinematics of the final-state nucleon. As a result, elastic FSI is included. The FSI from the NEUT cascade does not pose a double-counting problem [39].
Recently, a dedicated nuclear deexcitation event generator NucDeEx [3, 4, 6] was implemented in NEUT. The branching ratios of high-energy continuum excited states above the separation energy are calculated with the Hauser-Feshbach model [29], as implemented in the nuclear reaction code TALYS [34]. On the other hand, the gamma-ray branching ratios from low-lying discrete excited states of $^{11}\text{B}^{*}$ and $^{11}\text{C}^{*}$ are taken from experimental measurements [40]. It should be noted that NucDeEx enables to simulate deexcitation from various nuclei other than single nucleon knockout, $^{11}\text{C}^{*}$ in this analysis. A detailed description of the implementation in NEUT is given in Ref. [4].
@@ III Results and discussion
where $E_{\nu}(235.5\,\text{MeV})$ is the K-DAR neutrino energy, $m_{\mu}(105.7\text{ MeV})$ is the muon mass and $E_{\text{vis}}$ is the visible energy inside the detector. This latter quantity has contributions from the muon energy, the kinetic energy of protons, as well as from de-excitation photons. The event selection requires a muon in the final state and that the total visible energy satisfies $20\,\text{MeV}<E_{\text{vis}}<150\,\text{MeV}$.
@@
where $\sum T_{p}$ is the sum of proton kinetic energies and $E_{\gamma}$ is the energy of the gamma rays emitted by nuclear deexcitation of the residual nuclei. To match the signal definition as closely as possible, a muon in the final state is required and the total kinetic energy of the muon and proton(s), $T=T_{\mu}+\sum T_{p}$, must satisfy $20\,\text{MeV}<T<150\,\text{MeV}$.
@@
Separate covariance matrices are available for the statistical uncertainty, proton Birks’ constant uncertainty, and generator systematic uncertainties. During the calculation of the $\chi^{2}$, a selection is applied to the $E_{\text{vis}}$ values such that $45\,\text{MeV}\leq E_{\text{vis}}<120\,\text{MeV}$, in order to exclude bins with null cross section. The selection is applied to the events and covariance matrices before the $\chi^{2}$ calculation. The $E_{\text{vis}}$ selection corresponds to $4.8\,\text{MeV}\leq E_{m}<84.8\,\text{MeV}$. The lower bound of this $E_{m}$ selection is below the 1NKO. Taking this threshold into account, the $E_{\text{vis}}$ selection becomes $18.7\,\text{MeV}\leq E_{\text{vis}}<84.8\,\text{MeV}$, and three degrees of freedom from the covariance matrices and $\chi^{2}$ calculation are removed. Importantly, one of these bins (likely corresponding to the $1p_{3/2}$ peak) contains the 1NKO threshold itself, so the removal of this bin also removes the $1p_{3/2}$ peak from the $chi^{2}$ calculation.
@@
In an effort to isolate the effect of the NEUT FSI cascade from the modeling of the ground state and the de-excitation routine, each component is turned on sequentially and compared to the JSNS2 measurement in Figure 2. The obtained $\chi^{2}$ values, with and without the 1NKO threshold and the corresponding $p$-values, are shown in Table 2. To better understand the contribution of each individual bin to the overall $\chi^{2}$ value, a $\chi^{2}_{N-1}$ analysis is shown in Figures 3 and 4 for the cases without and with the 1NKO threshold respectively. In such an analysis, each bin is removed from the total $\chi^{2}$ calculation to isolate its effect. It should be noted that the bins around the $1p_{3/2}$ peak present large correlations, which can significantly alter the value of the $\chi^{2}$, and that the inclusion of the 1NKO threshold removes the $1p_{3/2}$ peak from the $\chi^{2}$ calculation entirely.
@@
The models in NEUT, without the NEUT cascade and NucDeEx deexcitation, are compared to the measurement in the left side of Figure 2. The SF and ED-RMF models both overestimate the ground-state $1p_{3/2}$ peak, whereas the SF model reproduces the height of the peak bin adequately. Before the $1p_{3/2}$ peak, the SF model has no strength, unlike the SF and ED-RMF models, as expected of the Dirac delta description of the ground-state in the SF model. All models struggle to reproduce the bin after the $1p_{3/2}$ peak, but in different ways – the SF and ED-RMF models overestimate the contribution, while the SF model significantly underestimates it. From Figure 3, only the bins adjacent to the ground-state peak contribute most significantly to the overall $\chi^{2}$ for the SF model. For the ED-RMF and SF models, both the $1p_{3/2}$ peak bin and its adjacent bins contribute most significantly. These observations follow the qualitative trends that can be expected from the removal energy distributions in Figure 1. When applying the 1NKO threshold, this is no longer the case for the SF and ED-RMF models, where the contribution to the $chi^{2}$ predominantly comes from the tail of the distribution. For the SF model the dominant $\chi^{2}$ contribution now comes from the bin just after the $1p_{3/2}$ peak. The second $1s_{1/2}$ peak around 40 MeV is inadequately described by all models. The lack of inelastic FSI means that the strength of nucleon rescatters that increases the size of the tail of the distribution is not present, as is expected from all three models. The $\chi^{2}$ values show that all models are rejected in the case where the NEUT cascade and NucDeEx are switched off.
@@
In the discussion so far, we have selected ED-RMF as the RMF-based model. Other nucleon potential models, such as RPWIA, EDAI, and rEDAI, lead to only slight changes in the results (see Appendix A). Furthermore, in the ED-RMF model, it is possible to vary the shell occupancies, as well as central values and widths of the Gaussian functions, as shown in Table. 1. The impact of varying this parameterisation is discussed in Appendix B.

### v5

@@ I Introduction
Precise modeling of neutrino–nucleus interactions is essential for current and future neutrino oscillation experiments, where systematic uncertainties associated with nuclear effects directly impact the interpretation of measured signals [9]. Current and future generation neutrino oscillation experiments heavily rely on Monte Carlo event generators to predict neutrino interactions in the GeV regime. Until recently, neutrino generators mostly relied on inclusive theoretical model implementations that predict differential cross section predictions only as a function of final-state lepton kinematics, even though the generators produce fully exclusive predictions for all particles in the interaction. Over the last few years, generators have undergone an overhaul that accommodates the implementation of more exclusive models by relying on the factorization approach [33, 10]. This has notably allowed for the implementation of shell-based models such as the Benhar Spectral Function (SF) [15, 17, 12]. Recently, the unfactorized energy-dependent relativistic mean field (ED-RMF) model was implemented into the NEUT generator using a tabulated hadron tensor approach [44]. These models have been benchmarked against exclusive electron scattering measurements, and their validation using exclusive neutrino scattering measurements has recently begun [28, 27, 25, 48].
@@
A major challenge in measuring neutrino scattering cross sections arises from the broadband nature of accelerator neutrino beams. The wide energy spectrum complicates the reconstruction of the incoming neutrino energy and smears out final-state observables, making it difficult to disentangle the modeling of the nuclear ground state from nuclear effects. An exception to this comes from neutrinos produced by kaon decay at rest (KDAR). In this process, a kaon undergoes a two-body decay producing a neutrino ($K^{+}\rightarrow\mu^{+}\nu_{\mu}$) with a well defined energy of 235.5 MeV, with a branching ratio of 63.6% [31].
@@
Recently, the JSNS2 collaboration has released its first measurement of the missing energy spectrum using neutrinos from a KDAR source on a carbon target [43]. This measurement can be used to study the neutron spectral function of carbon, which is a widely used nuclear target in experiments (for example, by experiments such as T2K [1], MINERvA [8], and NOvA [14]). These experiments operate at energies of $\approx 1$ GeV, so KDAR neutrinos are comparable to the regions of interest for accelerator-based experiments (such as T2K and NOvA, and in the future, DUNE [7] and Hyper-K [2]). Moreover, these energies are particularly relevant for testing the approximations on which neutrino interaction generators are built. For example, many generators rely on the plane wave impulse approximation (PWIA), which assumes that the neutrino scatters off a single isolated nucleon and that the outgoing particles propagate as undistorted plane waves, neglecting correlations and interactions with the residual nucleus. The energy region of the KDAR neutrinos is low enough that the energy transfers involved in charged-current (CC) interactions become sensitive to nuclear medium effects (for example, related to the nuclear ground state, Pauli blocking, or nucleon-nucleon correlations), which provides a unique opportunity to probe them directly.
@@
In this work, we focus on benchmarking three exclusive shell-based models implemented in the NEUT event generator [33] (version 6.0.3) against the JSNS2 measurement of KDAR neutrinos. We choose to focus on the NEUT event generator because, at the time of writing, it offers the widest choice of exclusive shell-based models within a single generator, allowing us to test the models in a consistent framework.
@@ II Interaction Models
$\tilde{E}=E_{k}-E_{k^{\prime}}-E_{p^{\prime}}+M$ (1)
where $E_{k}$, $E_{k^{\prime}}$, and $E_{p^{\prime}}$ represent the energy of the incoming lepton, outgoing lepton, and outgoing nucleon, respectively, while $M$ represents the target nucleon mass. Note that the removal energy is identical to the missing energy discussed here in the absence of final-state interactions (FSI). Recently, a new carbon SF with high removal energy resolution [12] has also been implemented in NEUT. This new SF (denoted as SF in this work) distinguishes between the ground state and the first and second excited states of $^{11}\text{B}^{*}$ (with excitation energies at 2.13 MeV and 5.02 MeV, as shown in Fig. 1) that were smeared out by coarse binning in the previous SF. Thus, it is expected to provide a more precise description of gamma-ray emissions by the nuclear deexcitation. Since the SF tables are obtained from exclusive electron scattering measurements $(e,e^{\prime}p)$, they provide only the spectral function for protons. To obtain the neutron spectral function, an additional correction is required. While assuming that the momentum distributions of protons and neutrons in symmetric nuclei are the same, a shift to the removal energy distribution is applied to neutrons, following the discussion in Refs. [13, 11].
@@
where $\Delta\tilde{E}=2.76$ MeV for the carbon nucleus [23]. This correction takes into account the fact that neutrons are more deeply bound in the nucleus than protons due to the absence of the Coulomb potential. In the case of SF, this correction approximately reproduces the removal energy corresponding to the first and second discrete excited states of $^{11}\text{C}^{*}$ (with excitation energies at 2.00 MeV and 4.80 MeV). The results of applying this correction for both SF and SF will be discussed in this paper. Pauli blocking is implemented using a simple step function, which requires the momentum of the outgoing nucleon to exceed the Fermi surface $p_{F}=209$ MeV [17]. In the high-removal energy and high-momentum region, the SF and SF models include a non-negligible contribution from short-range correlations (SRCs). In NEUT, events with $\tilde{E}>100\,$MeV or $|\mathbf{p}|>300$ MeV are defined as originating from SRCs. The SRCs, according to the NEUT definition, account for approximately 10% of the total integrated SF probability. Their contribution to the CCQE cross section is energy dependent: at the KDAR neutrino energy of 235.5 MeV it is about 1%, while at higher neutrino energies it can reach approximately 7%. For these events, the target nucleon is assumed to exist as part of a correlated nucleon pair in a back-to-back configuration, and the correlated partner nucleon has momentum opposite to that of the initial target nucleon. In such events, both nucleons undergo subsequent FSI.
Figure 1: Energy levels of 12C, 11B, and 11C relative to the 12C ground state based on Refs. [35, 37]. The excited states with $J^{\pi}=5/2^{-}$ of $^{11}\mathrm{C}$ and $^{11}\mathrm{B}$, shown in red, are not included in the SF [12] while the other three levels are included.
@@
where $m_{j}$ and $s$ are the third component of the total angular momentum of the hole state and the spin of the scattered nucleon state, respectively. $j$ denotes the total angular momentum, and $2j+1$ is the total occupancy of a given shell. The hadron current $J^{\mu}_{\kappa,m_{j},s}$ is described by single-particle RMF states [28] as
@@
Here, $\Psi_{\kappa}^{m_{j}}(\mathbf{p})$ is the initial-state RMF wavefunction, $\Gamma^{\mu}$ is the transition operator for the CCQE interaction in the CC2 form [39], and $\bar{\psi}_{s}(\mathbf{p}+\mathbf{q},\mathbf{p_{N}})$ is the final-state wavefunction that is the solution to the Dirac equation in the presence of a nuclear potential. $\mathbf{p}$ is the momentum of the initial nucleon, $\mathbf{p_{N}}$ is the momentum of the final-state nucleon, and $\mathbf{q}$ is the three momentum transfer. The model has multiple nuclear potentials available: ED-RMF [28], energy-dependent $A$-independent relativistic optical potentials (EDAI) [21], the real-only component of the EDAI optical potential (rEDAI) and no potential which simplifies to the relativistic plane wave impulse approximation (RPWIA). The inclusion of the nuclear potential in the final state is not present in the SF models and provides additional information about elastic FSI that the scattered nucleon undergoes. In order to model the nucleus in a similar fashion to the SF model, without the need for factorization, the missing energy profile, which is factorized from the precomputed hadron tensor tables, is modeled using Gaussian nuclear shells [44]; the Gaussian shell structure for carbon is given in Table 1 and a “background” shell contribution is included that is modeled in three distinct regions in $E_{m}$. The shell positions are chosen to be solutions to RMF equations or are taken from recent measurements where they exist. The shell occupancies and widths are chosen to mimic the SF [15, 16] following the procedure described in Ref. [25]. The ED-RMF potential has a natural implementation of Pauli blocking due to the orthogonality of the wavefunctions for the initial and final states [28].
@@
The NEUT FSI cascade is based on the mean free path of a particle travelling through the nuclear medium [33]. Particles are tracked and stepped by a fixed size of 0.2 fm through the nuclear medium until it leaves the nucleus. During this process, nucleon rescattering based on the Bertini model [19] and pion absorption based on Ref. [51] can occur, leading creation and ejection of other particles from the nucleus. The NEUT FSI cascade is inherently an inelastic process; during nucleon rescatters, both nucleons involved in the interaction are tracked through further reinteractions in the cascade. In contrast, the ED-RMF model incorporates the final-state potential into calculating the kinematics of the final-state nucleon. As a result, elastic FSI is included. The FSI from the NEUT cascade does not pose a double-counting problem [48].
Recently, a dedicated nuclear deexcitation event generator NucDeEx [3, 4, 6] was implemented in NEUT. The branching ratios of high-energy continuum excited states above the separation energy are calculated with the Hauser-Feshbach model [32], as implemented in the nuclear reaction code TALYS [41]. NucDeEx simulates emission of various light particles, such as nucleons, deuterons, tritons, 3He, and alpha particles. On the other hand, the gamma-ray branching ratios from low-lying discrete excited states of $^{11}\text{B}^{*}$ and $^{11}\text{C}^{*}$ are taken from experimental measurements [49]. It should be noted that NucDeEx enables to simulate deexcitation from various nuclei other than single nucleon knockout, $^{11}\text{C}^{*}$ in this analysis. A detailed description of the implementation in NEUT is given in Ref. [4].
@@ III Results and discussion
where $E_{\nu}(235.5\,\text{MeV})$ is the KDAR neutrino energy, $m_{\mu}(105.7\text{ MeV})$ is the muon mass and $E_{\text{vis}}$ is the visible energy inside the detector. This latter quantity has contributions from the muon energy, the kinetic energy of protons, as well as from de-excitation photons. The event selection requires a muon in the final state and that the total visible energy satisfies $20\,\text{MeV}<E_{\text{vis}}<150\,\text{MeV}$.
@@
where $\sum T_{p}$ is the sum of proton kinetic energies and $E_{\gamma}$ is the energy of the gamma rays emitted by nuclear deexcitation of the residual nuclei. To match the signal definition as closely as possible, a muon in the final state is required and the total kinetic energy of the muon and proton(s), $T=T_{\mu}+\sum T_{p}$, must satisfy $20\,\text{MeV}<T<150\,\text{MeV}$. It should be noted that the kinetic energy of other charged particles, such as deuterons and alpha particles, is not included in this missing energy calculation. Although NucDeEx can simulate the emission of these particles, their visible energy is expected to be much lower than their kinetic energy due to stronger quenching effect than that of protons. Since the quenching effect for these heavy charged particles is not discussed in Ref. [43], we did not include their contribution in the missing energy calculation.
@@
Separate covariance matrices are available for the statistical uncertainty, proton Birks’ constant uncertainty, and generator systematic uncertainties. During the calculation of the $\chi^{2}$, a selection is applied to the $E_{\text{vis}}$ values such that $45\,\text{MeV}\leq E_{\text{vis}}<120\,\text{MeV}$, in order to exclude bins with null cross section. The selection is applied to the events and covariance matrices before the $\chi^{2}$ calculation. The $E_{\text{vis}}$ selection corresponds to $4.8\,\text{MeV}\leq E_{m}<84.8\,\text{MeV}$. The lower bound of this $E_{m}$ selection is below the 1NKO.
Computing the $\chi^{2}$ and corresponding $p$-values in the case where the 1NKO is applied requires the introduction of two modifications to the measurement and associated covariance matrix. Taking this threshold into account, the $E_{\text{vis}}$ selection becomes $18.7\,\text{MeV}\leq E_{\text{vis}}<84.8\,\text{MeV}$, which removes three degrees of freedom from the $\chi^{2}$ calculation. The corresponding rows and columns from the covariance matrices are also removed. This operation is justified under the assumption of a gaussian distribution of errors, which is inherent to the treatment via a covariance matrix. Since this measurement is presented in the unfolded truth space, it is assumed that possible spill-over artifacts from adjacent bins were corrected for during the unfolding procedure and the binning size was chosen accordingly. Importantly, one of the bins which is removed (likely corresponding to the $1p_{3/2}$ peak) contains the 1NKO threshold itself, so the removal of this bin also removes the $1p_{3/2}$ peak from the $\chi^{2}$ calculation. The bins around the $1p_{3/2}$ peak present large correlations, which can significantly alter the value of the $\chi^{2}$.
@@
In an effort to isolate the effect of the NEUT FSI cascade from the modeling of the ground state and the de-excitation routine, each component is turned on sequentially and compared to the JSNS2 measurement in Figure 3. The obtained $\chi^{2}$ values, with and without the 1NKO threshold and the corresponding $p$-values, are shown in Table 2. To better understand the contribution of each individual bin to the overall $\chi^{2}$ value, a $\chi^{2}_{N-1}$ analysis is shown in Figures 4 and 5 for the cases without and with the 1NKO threshold respectively. In such an analysis, each bin, along with its corresponding covariance matrix row and column, is removed from the total $\chi^{2}$ calculation to isolate its effect (but still accounting for the remaining correlations).
@@
The models in NEUT, without the NEUT cascade and NucDeEx deexcitation, are compared to the measurement in the left side of Figure 3. The SF and ED-RMF models both overestimate the ground-state $1p_{3/2}$ peak, whereas the SF model reproduces the height of the peak bin adequately. Before the $1p_{3/2}$ peak, the SF model has no strength, unlike the SF and ED-RMF models, as expected of the Dirac delta description of the ground-state in the SF model. All models struggle to reproduce the bin after the $1p_{3/2}$ peak, but in different ways – the SF and ED-RMF models overestimate the contribution, while the SF model significantly underestimates it. From Figure 4, only the bins adjacent to the ground-state peak contribute most significantly to the overall $\chi^{2}$ for the SF model. For the ED-RMF and SF models, both the $1p_{3/2}$ peak bin and its adjacent bins contribute most significantly. These observations follow the qualitative trends that can be expected from the removal energy distributions in Figure 2. When applying the 1NKO threshold, this is no longer the case for the SF and ED-RMF models, where the contribution to the $\chi^{2}$ predominantly comes from the tail of the distribution. For the SF model the dominant $\chi^{2}$ contribution now comes from the bin just after the $1p_{3/2}$ peak. The second $1s_{1/2}$ peak around 40 MeV is inadequately described by all models. The lack of inelastic FSI means that the strength of nucleon rescatters that increases the size of the tail of the distribution is not present, as is expected from all three models. The $\chi^{2}$ values show that all models are rejected in the case where the NEUT cascade and NucDeEx are switched off.
@@
In the discussion so far, ED-RMF is used as the default RMF-based model. Other nucleon potential models, such as RPWIA, EDAI, and rEDAI, lead to only slight changes in the results (see Appendix A). Furthermore, in the ED-RMF model, it is possible to vary the shell occupancies, as well as central values and widths of the Gaussian functions, as shown in Table. 1. The impact of varying this parameterization is discussed in Appendix B.
Additionally, the impact of varying the FSI strength in the NEUT cascade is shown in Appendix C. While this work uses the NEUT cascade, investigating other generator cascades, such as NuWro [26] as shown in Ref. [24], is necessary to draw more general conclusions about how FSI affects agreement with the measurement.
[…]